Solve the following system of simultaneous equations using the elimination method.
x + 8y = 4
5x + 2y = 96

Answer by oscargut(2103)

x + 8y = 4
5x + 2y = 96
Multiply the 2nd eq by -4
x+8y =4
-20x-8y = -384
adding:
-19x= -380
x = 20
y = -2
I can help you with more questions at: mthman@gmail.com
Thanks